OpenStudy (anonymous):

a card is drawn froma well shuffled deck of 52 cards. find the probability that the card is a black card or a 8

OpenStudy (anonymous):

total black cards = 26 total 8 cards = 4 but two of these are also included in the black cards Thus, Total favorable cases = 26 + 4 -2 = 28 \[Probability =\frac{ FavourableCases }{AllCases } \] P = 28/ 52

OpenStudy (anonymous):

thank you so much for your time and effort i appreciate it!

OpenStudy (anonymous):

Its been a pleasure :)